# Distributed Constrained Online Nonconvex Optimization with Compressed Communication

### v1

@@ V Conclusions
This paper studied the distributed online nonconvex optimization problem with time-varying constraints. To better utilize communication resources, we proposed a distributed online primal–dual algorithm with compressed communication. More importantly, the algorithm was able to handle time-varying communication topologies. We showed that the algorithm established sublinear network regret and cumulative constraint violation bounds. Moreover, the network cumulative constraint violation bounds were further reduced when Slater’s condition held. These bounds were comparable to the state-of-the-art results established by existing distributed online algorithms with perfect communication, even in the context of distributed online convex optimization with (time-varying) inequality constraints. In the future, we plan to investigate distributed bandit nonconvex optimization with time-varying constraints since gradient information is unavailable in many real-world applications.
@@ A. Useful Lemmas
$={{\tilde{p}}^{2}}{\mathbf{E}_{\mathcal{C}}}[{\|{{z_{i,t+1}}-{\hat{z}_{i,t}}-{s_{t+1}}\mathcal{C}({({{z_{i,t+1}}-{\hat{z}_{i,t}}})/{s_{t+1}}})}\|_{d}^{2}}]$
@@
${{\tilde{\vartheta}}_{3}}=\frac{{16n{\tau^{2}}}}{{{\lambda^{2}}({1-{\lambda^{2}}})}}{\Big{(}{\sum\limits_{i=1}^{n}{\|{{\hat{z}_{i,1}}}\|}}\Big{)}^{2}},{{\tilde{\vartheta}}_{4}}=\frac{{16{n^{2}}{\tau^{2}}}}{{{{\left({1-\lambda}\right)}^{2}}}}+32.$
@@
$\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\langle{\nabla{f_{t}}({{x_{i,t}}}),{x_{i,t}}-y}\rangle}]}}\leq 2nR(\mathbb{X})L{\vartheta_{1}}+{\vartheta_{2}}\sum\limits_{t=1}^{T}{{\alpha_{t}}}+2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}\sum\limits_{t=1}^{T}{{s_{t}}}$
@@
${\vartheta_{2}}=2G_{1}^{2}+4n^{2}R{(\mathbb{X})^{2}}{L^{2}}\tilde{\vartheta}_{2}^{2},{\vartheta_{3}}=2G_{2}^{2}{{\tilde{\vartheta}}_{3}},{\vartheta_{4}}=\frac{{4\max\{{1,2G_{2}^{2}{{\tilde{\vartheta}}_{4}}}\}}}{{\min\{{1,\frac{1}{{2{\gamma_{0}}}}}\}}},{\vartheta_{5}}={n{G_{1}}{G_{2}}{{\tilde{\vartheta}}_{2}}},$
${\vartheta_{6}}={1+G_{2}^{2}{\gamma_{0}}{{\tilde{\vartheta}}_{2}}}.$
@@
$\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\langle{\nabla{f_{i,t}}({{x_{i,t}}}),{x_{i,t}}-y}\rangle}]}}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{L\|{{x_{i,t}}-{x_{j,t}}}\|\|{{x_{i,t}}-y}\|}]}}}$
$\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\langle{\nabla{f_{i,t}}({{x_{i,t}}}),{x_{i,t}}-y}\rangle}]}}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{2R(\mathbb{X})L\|{{x_{i,t}}-{x_{j,t}}}\|}]}}},\forall y\in\mathbb{X},$ (65)
where last inequality holds due to (3).
For the second term on the right-hand side of (65), from (46c), we have
$\;\;\;\;\;\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{2R(\mathbb{X})L\|{{x_{i,t}}-{x_{j,t}}}\|}]}}}$
$\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{2R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}\|{\varepsilon_{i,t}^{z}}\|}]}}$
$\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{2R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}\|{{z_{i,t+1}}-{{\hat{z}}_{i,t+1}}}\|}]}}+\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{2R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}\|{{x_{i,t}}-{z_{i,t+1}}}\|}]}}$
$\leq 2nR(\mathbb{X})L{\vartheta_{1}}+2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}\sum\limits_{t=1}^{T}{{s_{t}}}+\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}\Big{[}{4nR{{(\mathbb{X})}^{2}}{L^{2}}\tilde{\vartheta}_{2}^{2}{\alpha_{t}}+\frac{{{{\|{{x_{i,t}}-{z_{i,t+1}}}\|}^{2}}}}{{4n{\alpha_{t}}}}}\Big{]}}}$
$\leq 2nR(\mathbb{X})L{\vartheta_{1}}+2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}\sum\limits_{t=1}^{T}{{s_{t}}}+4n^{2}R{(\mathbb{X})^{2}}{L^{2}}\tilde{\vartheta}_{2}^{2}\sum\limits_{t=1}^{T}{{\alpha_{t}}}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\frac{{{\mathbf{E}_{\mathcal{C}}}[{{{\|{{x_{i,t}}-{z_{i,t+1}}}\|}^{2}}}]}}{{4{\alpha_{t}}}}}},$ (66)
where the second inequality holds due to $\varepsilon_{i,t}^{z}={{\hat{z}}_{i,t+1}}-{x_{i,t}}$; and the third inequality holds since (45) holds and $\{{{s_{t}}}\}$ is nonincreasing.
@@
where the second inequality holds due to the nonexpansive property of the projection ${[\cdot]_{+}}$; and the last inequality holds due to (5).
@@
where the first inequality holds due to ${g_{t}}(x)={\rm{col}}({{g_{1,t}}(x),\cdot\cdot\cdot,{g_{n,t}}(x)})$; the second inequality holds due to the nonexpansive property of the projection ${[\cdot]_{+}}$; and the last inequality holds due to (5).
@@ B. Proof of Theorem 1
[…]
@@
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+\frac{{2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}{s_{0}}}}{{1-{\theta_{2}}}}{T^{1-{\theta_{2}}}}$
@@
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+4n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}{s_{0}}\log(T)$
@@
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}{\rm Z_{1}}{s_{0}}$
@@
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}{\rm Z_{2}}{s_{0}}$
@@
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}{\rm Z_{3}}{s_{0}}$
@@ C. Proof of Theorem 2
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\frac{{2{\vartheta_{2}}{\alpha_{0}}\mu}}{{1-\mu}}\sqrt{T}+\frac{{2n\tilde{p}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}{s_{0}}\mu}}{{1-\mu}}$

### v2

@@ V Conclusions
This paper studied the distributed online nonconvex optimization problem with time-varying constraints. To better utilize communication resources, we proposed a distributed online primal–dual algorithm with compressed communication. More importantly, the algorithm was able to handle time-varying communication topologies. We showed that the algorithm established sublinear network regret and cumulative constraint violation bounds. Moreover, the network cumulative constraint violation bounds were further reduced when Slater’s condition held. In the future, we plan to investigate distributed bandit nonconvex optimization with time-varying constraints since gradient information is unavailable in many real-world applications.
@@ A. Useful Lemmas
$\leq{{\tilde{p}}^{2}}{\mathbf{E}_{\mathcal{C}}}[{\|{{z_{i,t+1}}-{\hat{z}_{i,t}}-{s_{t+1}}\mathcal{C}({({{z_{i,t+1}}-{\hat{z}_{i,t}}})/{s_{t+1}}})}\|_{d}^{2}}]$
@@
${{\tilde{\vartheta}}_{3}}=\frac{{16n{\tau^{2}}}}{{{\lambda^{2}}({1-{\lambda^{2}}})}}{\Big{(}{\sum\limits_{i=1}^{n}{\|{{\hat{z}_{i,1}}}\|}}\Big{)}^{2}},{{\tilde{\vartheta}}_{4}}=\frac{{16{n^{2}}{\tau^{2}}}}{{{{({1-\lambda})}^{2}}}}+32.$
@@
$\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\langle{\nabla{f_{t}}({{x_{i,t}}}),{x_{i,t}}-y}\rangle}]}}\leq\hat{\vartheta}+{\vartheta_{2}}\sum\limits_{t=1}^{T}{{\alpha_{t}}}+\tilde{p}\sqrt{C}{{\tilde{\vartheta}}_{1}}\sum\limits_{t=1}^{T}{{s_{t}}}$
@@
$\hat{\vartheta}=2R(\mathbb{X})L{\vartheta_{1}}+{G_{1}}{\vartheta_{1}},{{\tilde{\vartheta}}_{1}}=2R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}+{G_{1}}{{\tilde{\vartheta}}_{2}},{\vartheta_{2}}=2G_{1}^{2}+\tilde{\vartheta}_{1}^{2},{\vartheta_{3}}=2G_{2}^{2}{{\tilde{\vartheta}}_{3}},$
${\vartheta_{4}}=\frac{{4\max\{{1,2G_{2}^{2}{{\tilde{\vartheta}}_{4}}}\}}}{{\min\{{1,\frac{1}{{2{\gamma_{0}}}}}\}}},{\vartheta_{5}}={n{G_{1}}{G_{2}}{{\tilde{\vartheta}}_{2}}},{\vartheta_{6}}={1+G_{2}^{2}{\gamma_{0}}{{\tilde{\vartheta}}_{2}}}.$
@@
$\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\langle{\nabla{f_{t}}({{x_{i,t}}}),{{\bar{z}}_{t}}-y}\rangle}]}}+\frac{1}{{{n^{2}}}}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{\nabla{f_{j,t}}({{x_{i,t}}})}\|\|{{x_{i,t}}-{{\bar{z}}_{t}}}\|}]}}}$
$\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\langle{\nabla{f_{i,t}}({{x_{i,t}}}),{{\bar{z}}_{t}}-y}\rangle}]}}+\frac{{2R(\mathbb{X})L}}{{{n^{2}}}}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{x_{i,t}}-{x_{j,t}}}\|}]}}}$
$\;\;+\frac{{{G_{1}}}}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{x_{i,t}}-{{\bar{z}}_{t}}}\|}]}}$
$\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\langle{\nabla{f_{i,t}}({{x_{i,t}}}),{x_{i,t}}-y}\rangle}]}}+\frac{{2R(\mathbb{X})L}}{{{n^{2}}}}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{x_{i,t}}-{x_{j,t}}}\|}]}}}$
$\;\;+\frac{{2{G_{1}}}}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{x_{i,t}}-{{\bar{z}}_{t}}}\|}]}}.$ (61)
From (42c), we have
$\frac{1}{n^{2}}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{2R(\mathbb{X})L\|{{x_{i,t}}-{x_{j,t}}}\|}]}}}\leq 2R(\mathbb{X})L{\vartheta_{1}}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{2R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}\|{\varepsilon_{i,t}^{z}}\|}]}}.$ (62)
From (54), we have
$\frac{{2{G_{1}}}}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{\bar{z}}_{i,t}}-{{\bar{z}}_{t}}}\|}]}}\leq{G_{1}}{\vartheta_{1}}+\frac{{{G_{1}}{{\tilde{\vartheta}}_{2}}}}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\|{\varepsilon_{i,t}^{z}}\|}]}}.$ (63)
From (41), we have
$\;\;\;\;\;\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{{{\tilde{\vartheta}}_{1}}\|{\varepsilon_{i,t}^{z}}\|}]}}$
$=\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{{{\tilde{\vartheta}}_{1}}\|{{z_{i,t+1}}-{{\hat{z}}_{i,t+1}}}\|}]}}+\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{{{\tilde{\vartheta}}_{1}}\|{{z_{i,t+1}}-{x_{i,t}}}\|}]}}$
$\leq\tilde{p}\sqrt{C}{{\tilde{\vartheta}}_{1}}\sum\limits_{t=1}^{T}{{s_{t+1}}}+\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}\Big{[}{\tilde{\vartheta}_{1}^{2}{\alpha_{t}}+\frac{{{{\|{{z_{i,t+1}}-{x_{i,t}}}\|}^{2}}}}{{4{\alpha_{t}}}}}\Big{]}}}.$ (64)
@@
where the second inequality holds due to the nonexpansive property of the projection $\cdot$$_{+}$}; and the last inequality holds due to (5).
@@
where the first inequality holds due to ${g_{t}}(x)={\rm{col}}({{g_{1,t}}(x),\cdot\cdot\cdot,{g_{n,t}}(x)})$; the second inequality holds due to the nonexpansive property of the projection $\cdot$$_{+}$}; and the last inequality holds due to (5).
@@ B. Proof of Theorem 1
$\sum\limits_{t=1}^{T}{{\alpha_{t}}}={\alpha_{0}}\sum\limits_{t=1}^{T}{\frac{1}{{{t^{{\theta_{1}}}}}}}={\alpha_{0}}\Big{(}\sum\limits_{t=2}^{T}{\frac{1}{{{t^{{\theta_{1}}}}}}}+1\Big{)}\leq{\alpha_{0}}\Big{(}\int_{1}^{T}{\frac{1}{{{t^{{\theta_{1}}}}}}dt}+1\Big{)}\leq\frac{{{{\alpha_{0}}T^{1-{\theta_{1}}}}}}{{1-{\theta_{1}}}}.$ (73)
@@
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq{\hat{\vartheta}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+\frac{{\tilde{p}\sqrt{C}{{\tilde{\vartheta}}_{1}}{s_{0}}}}{{1-{\theta_{2}}}}{T^{1-{\theta_{2}}}}$
@@
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq{\hat{\vartheta}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+2\tilde{p}\sqrt{C}{{\tilde{\vartheta}}_{1}}{s_{0}}\log(T)$
@@
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq{\hat{\vartheta}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+\tilde{p}\sqrt{C}{{\tilde{\vartheta}}_{1}}{\rm Z_{1}}{s_{0}}$
@@
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq{\hat{\vartheta}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+\tilde{p}\sqrt{C}{{\tilde{\vartheta}}_{1}}{\rm Z_{2}}{s_{0}}$
@@
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq{\hat{\vartheta}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+\tilde{p}\sqrt{C}{{\tilde{\vartheta}}_{1}}{\rm Z_{3}}{s_{0}}$
@@ C. Proof of Theorem 2
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq{\hat{\vartheta}}+\frac{{2{\vartheta_{2}}{\alpha_{0}}\mu}}{{1-\mu}}\sqrt{T}+\frac{{\tilde{p}\sqrt{C}{{\tilde{\vartheta}}_{1}}{s_{0}}\mu}}{{1-\mu}}$